OpenStudy (anonymous):

A student writes an incorrect step while checking if the sum of the measures of the two remote interior angles of triangle ABC below is equal to the measure of the exterior angle. A triangle ABC is shown. The base of the triangle extends into a straight line. The angle formed between this straight line and the edge of the triangle is marked as p. The angle adjacent to w is marked as o, and the other two angles inside the triangle are marked as m and n. Step 1: m∠m + m∠n + m∠o = 180 degrees (sum of angles of a triangle) Step 2: m∠p − m∠o = 90 degrees (alternate interior angles)

OpenStudy (anonymous):

In which step did the student first make a mistake and how can it be corrected? Step 1; it should be m∠m + m∠n + m∠o = 90 degrees (corresponding angles) Step 1; it should be m∠m + m∠n + m∠o = 90 degrees (adjacent angles) Step 2; it should be m∠o + m∠p = 180 degrees (alternate exterior angles) Step 2; it should be m∠o + m∠p = 180 degrees (supplementary angle

OpenStudy (anonymous):

I used to know this stuff. Not anymore, sorry

OpenStudy (anonymous):

I assume you meant the angle adjacent to p is marked as o.  Always draw a picture if you can. Maybe you can tell from the picture what step is wrong and how it should be corrected?